# ReferDINO: Referring Video Object Segmentation with Visual Grounding Foundations

### v1

@@ 4.3 Confidence-aware Query Pruning
Foundational visual grounding models typically utilize a large set of query embeddings to store extensive object information, e.g., GroundingDINO uses $N_{q}=900$ queries in the cross-modal decoder. Iteratively processing such a large amount of queries significantly limits the efficiency, especially in video processing. However, directly reducing these queries can compromise the well-pretrained object knowledge. To solve this dilemma, we propose a confidence-aware query pruning strategy to iteratively reject the low-confidence queries at each decoder layer, deriving only a compact set of important queries for subsequent computations, as shown in Figure 4.
@@
Specifically, the cross-modal decoder is stacked by 6 layers, and each layer consists of a self-attention, a cross-attention with image features and a cross-attention with text features. Let $\boldsymbol{Q}_{l}\in\mathbb{R}^{N_{l}\times d}$ denote the output query embeddings of the $l$-th decoder layer, where $N_{0}=N_{q}$. We reuse the attention weights in the decoder layer to compute a confidence score for each query as follows:
where $s_{j}$ is the confidence of $j$-th query, $\boldsymbol{A}^{s}\in\mathbb{R}^{N_{l}\times N_{l}}$ denotes the self-attention weight, $\boldsymbol{A}^{c}\in\mathbb{R}^{L\times N_{l}}$ denotes the transposed cross-attention weight with $L$ text tokens. The former term represents the average attention received by $j$-th query from other queries. A query receiving high attention from others typically indicates that it is irreplaceable. The latter maximum term measures the probability that $j$-th query (i.e., object) is mentioned in the text. The combination of these two terms represents the query importance and the relevance to text descriptions. Based on this score, we filter out $p$of the low-confidence object queries at each layer, finally yielding a compact set of $N_{s}$ object embeddings, where $N_{s}\ll N_{q}$. This strategy significantly reduces the computation costs without compromising performance, enabling our ReferDINO to satisfy the efficiency demands of video processing.
@@ 4.4 Training and Inference
ReferDINO ultimately produces $N_{s}$ object prediction sequences $\boldsymbol{p}=\{\boldsymbol{p}_{i}\}_{i=1}^{N_{s}}$ for a video-text pair, and each sequence is represented by $\boldsymbol{p}_{i}=\{\boldsymbol{c}^{t}_{i},\boldsymbol{b}^{t}_{i},\boldsymbol{m}^{t}_{i}\}_{t=1}^{T}$, which denotes the binary classification probability, bounding box and mask of the $i$-th object query on $t$-th frame. Note that there is only one object sequence correctly corresponding to the text description.
Training. Suppose the ground truth object sequence to be $\boldsymbol{y}=\{\boldsymbol{c}^{t},\boldsymbol{b}^{t},\boldsymbol{m}^{t}\}_{t=1}^{T}$. During training, we select the sequence with the lowest matching cost as the positive and assign the remaining sequences as negative. The matching cost is defined as follows:
The matching cost is computed on individual frames and normalized by the frame number. Here, $\mathcal{L}_{\text{cls}}$ is the focal loss that supervises the binary classification prediction. $\mathcal{L}_{\text{box}}$ sums up the L1 loss and GIoU loss. $\mathcal{L}_{\text{mask}}$ is the combination of DICE loss, binary mask focal loss and projection loss [31]. $\lambda_{\text{cls}}$, $\lambda_{\text{box}}$ and $\lambda_{\text{mask}}$ are scalar weights of individual losses. The model is optimized end-to-end by minimizing the total loss $\mathcal{L}_{\text{total}}$ for positive sequences and only the classification loss $\mathcal{L}_{\text{cls}}$ for negative sequences.
Inference. At inference, we select the best sequence with the highest average classification score, and its index is denoted as follows:
$\sigma=\underset{i\in[1,N_{s}]}{\arg\max}\frac{\sum_{t=1}^{T}\boldsymbol{c}^{t}_{i}}{T}.$ (6)
5 Experiments
5.1 Experimental Setup
Datasets. We evaluate ReferDINO on five public benchmarks: Ref-YouTube-VOS [30], Ref-DAVIS17 [15], A2D-Sentences, JHMDB-Sentences [6] and MeViS [5]. Ref-YouTube-VOS is a large-scale dataset covering 3,978 videos with 15K text descriptions. Ref-DAVIS17, A2D-Sentences and JHMDB-Sentences are created by annotating additional text descriptions on the original DAVIS17 [26], A2D [13] and JHMDB datasets, respectively. Ref-DAVIS17 contains 90 videos and 1.5K text descriptions. A2D-Sentences consists of 3.7K videos and 6.6K text descriptions. JHMDB-Sentences provides 928 videos, each with a description. MeViS is a recently established benchmark that poses new challenges in cross-modal motion understanding, including 2K videos and 28K text descriptions.
@@ 5.1 Experimental Setup
Grounded-SAM2 [1] arXiv’24 54.4 51.8 57.0 60.9 57.4 62.6 47.5 59.0 62.5 38.9 70.8 70.5
ReferDINO (ours) - 66.4 64.4 68.4 66.8 63.1 70.5 58.9 80.2 72.3 45.6 74.2 73.1
@@
Grounded-SAM2 [1] arXiv’24 64.8 62.5 67.0 66.2 62.6 69.7 54.7 67.8 68.5 42.5 72.1 72.1
ReferDINO (ours) - 69.3 67.0 71.5 68.9 65.1 72.9 61.1 82.1 73.6 46.6 74.2 73.2
Table 1: Comparison on Ref-YouTube-VOS, Ref-DAVIS17, A2D-Sentences and JHMDB-Sentences. †The results are obtained from [12].
@@
Metrics. Following previous works, we use region similarity $\mathcal{J}$ (average IoU), contour accuracy $\mathcal{F}$ (mean boundary similarity), and their average $\mathcal{J\&F}$ on MeViS, Ref-YouTube-VOS and Ref-DAVIS17. For A2D-Sentences and JHMDB-Sentences, we employ mAP, overall IoU (oIoU), and mean IoU (mIoU) metrics. All evaluations are conducted using the official code or online platforms.
@@
Protocols. We follow the experimental protocols established in previous works. Specifically, the model trained on the training set of Ref-YouTube-VOS is directly evaluated on the validation splits of Ref-YouTube-VOS and Ref-DAVIS17. Similarly, the model trained on A2D-Sentences is directly evaluated on A2D-Sentences and JHMDB-Sentences. The model is first pretrained on referring image segmentation datasets RefCOCO/+/g [14, 23], and then trained on the RVOS datasets, except for MeViS where the model is trained directly, following [5, 8].
@@
Implementation Details. Our model is built upon the pretrained GroundingDINO, which uses Swin Transformer as the image backbone and BERT as the text backbone. The official source releases two GroundingDINO checkpoints: Swin-T and Swin-B, both of which are covered in our experiments. We freeze the GroundingDINO’s backbones and finetune the cross-modal Transformer with LoRA techniques [10], where the rank is set to $32$. We set $\alpha=0.1$, $L_{t}=3$, and $L_{m}=3$. In the query pruning, we set the drop rate as $50\%$. For MeViS dataset involving multiple target objects, we follow the practice in [5, 8] to select multiple object trajectories with classification scores higher than a threshold $\sigma=0.2$. For the other datasets, we select the best object trajectory by Equation 6.
5.2 Main Results
Comparisons with SOTA Methods. As shown in Tables 1 and 2, our ReferDINO significantly and consistently outperforms SOTA methods across all five RVOS datasets. Specifically, with the Swin-T backbone, our ReferDINO achieves 48.0% $\mathcal{J}\&\mathcal{F}$ on the challenging dataset MeViS, outperforming existing SOTA by 1.6%. On the competitive dataset Ref-YouTube-VOS, our ReferDINO achieves 66.4% $\mathcal{J}\&\mathcal{F}$, improving SOTA performance by 2.8%. When a larger backbone Swin-B is applied, ReferDINO further improves $\mathcal{J}\&\mathcal{F}$ to 69.3% on Ref-YouTube-VOS, surpassing SOTA by more than 2.2%. Consistent performance improvements are observed across the other datasets, which demonstrates the superiority of our ReferDINO model.
Comparisons with GroundingDINO+SAM2. To demonstrate the effectiveness of our module designs upon GroundingDINO, we compare against a recent baseline Grounded-SAM2 [1], which uses GroundingDINO to identify objects in a single frame and then feed the box outputs into SAM2 [27] for segmenting throughout the video. We run with the official code and adopt the sam2_hiera_large version of SAM2. As shown in Tables 1 and  2, the performance of Grounded-SAM2 heavily depends on the detection quality of GroundingDINO. For example, switching GroundingDINO from Swin-B to Swin-T results in a drastic 10.4% drop in $\mathcal{J}\&\mathcal{F}$ on Ref-YouTube-VOS. Besides, such simple model ensemble cannot effectively address the RVOS task. Across various datasets and backbones, our ReferDINO consistently outperforms Grounded-SAM2 by a significant margin. This demonstrates the effectiveness of our approach in adapting GroundingDINO to RVOS task.
@@ 5.2 Main Results
Grounded-SAM2 [1] arXiv’24 37.4 31.0 43.7
ReferDINO (ours) - 48.0 43.6 52.3
@@
Grounded-SAM2 [1] arXiv’24 40.5 34.5 46.4
ReferDINO (ours) - 49.3 44.7 53.9
@@
Figure 5: Qualitative comparison of our ReferDINO with others.
@@ 5.3 Ablation Studies
Our model is built upon GroundingDINO with three elaborated designs, which we ablate in this section. We use MeViS as the main dataset as it is the largest.
[…]
Ours 48.0 43.6 52.3
@@
Table 3: Ablation study of the temporal aggregator.
@@
Dot-product 45.0 41.4 48.5
CondInst [31] 44.4 40.7 48.1
Ours 48.0 43.6 52.3
w/o CA 47.6 43.1 52.0
w/o DCA 45.3 41.8 48.7
Table 4: Ablation study of the mask decoder.
Temporal Enhancer. Our object-consistent temporal enhancer consists of a memory-augmented tracker and a cross-modal temporal decoder. We analyze their individual contributions in Table 3. First, while the object queries in GroundingDINO is naturally aligned across frames to some extent, adding a tracker can still improve the temporal consistency, leading to a 0.4% improvement in $\mathcal{J}\&\mathcal{F}$. Second, in the temporal decoder is essential for understanding temporal dynamics and object motion, providing 2.2% improvements in $\mathcal{J}\&\mathcal{F}$. These results demonstrates the effectiveness of our object-consistent temporal enhancer.

### v2

@@ 4.3 Confidence-aware Query Pruning
Foundational visual grounding models typically utilize a large set of query embeddings to store extensive object information, e.g., GroundingDINO uses $N_{q}=900$ queries. Iteratively processing such a large amount of queries significantly limits the efficiency, especially in video processing. However, directly reducing these queries can compromise the well-pretrained object knowledge. To solve this dilemma, we design a confidence-aware query pruning strategy to progressly identify and reject low-confidence queries at each decoder layer, as shown in Figure 4.
Specifically, the cross-modal decoder is stacked by $L$ layers, and each layer consists of a self-attention, a cross-attention with image features and a cross-attention with text features. Let $\boldsymbol{Q}_{l}\in\mathbb{R}^{N_{l}\times d}$ denote the output query embeddings of the $l$-th decoder layer, where $N_{0}=N_{q}$. We reuse the attention weights in the decoder layer to compute a confidence score for each query as follows:
where $s_{j}$ is the confidence of $j$-th query, $\boldsymbol{A}^{s}\in\mathbb{R}^{N_{l}\times N_{l}}$ denotes the self-attention weight, $\boldsymbol{A}^{c}\in\mathbb{R}^{K\times N_{l}}$ denotes the transposed cross-attention weight with $K$ text tokens. The former term represents the average attention received by $j$-th query from other queries. A query receiving high attention from others typically implies that it is irreplaceable. The latter maximum term measures the probability that $j$-th object query is mentioned in the text. The combination of these two terms comprehensively reflects the query importance. Based on this score, we retain only $1/k$ high-confidence object queries at each layer, finally yielding a compact set of $N_{s}$ object embeddings, where $N_{s}\ll N_{q}$. This strategy significantly reduces the computation costs while preserving the pretrianed object knowledge, enabling ReferDINO to achieve real-time efficiency.
Time Complexity Analysis. Let $N=N_{q}$, total time complexity of the original decoder is $O\left(L(N^{2}d+Nd^{2})\right)$. If we retain only $1/k$ queries per layer, the total query number would decrease exponentially, and the total time complexity becomes $O\left((\frac{k^{2}}{k^{2}-1})N^{2}d+(\frac{k}{k-1})Nd^{2}\right)$, independent of the decoder depth $L$. In practice, this computational improvement is significant. For GroundingDINO with $N=900$, $L=6$, $d=256$, even when $k=2$, the decoder computational cost can be reduced to 24.7%. More detailed derivation is presented in the supplementary A.
@@ 4.4 Training and Inference
ReferDINO ultimately produces $N_{s}$ object prediction sequences $\boldsymbol{p}=\{\boldsymbol{p}_{i}\}_{i=1}^{N_{s}}$ for a video-text pair, and each sequence is represented by $\boldsymbol{p}_{i}=\{\boldsymbol{c}^{t}_{i},\boldsymbol{b}^{t}_{i},\boldsymbol{m}^{t}_{i}\}_{t=1}^{T}$, which denotes the binary classification probability, bounding box and mask of the $i$-th object query on $t$-th frame.
@@
G-DINO+SH - 63.9 62.0 65.8 62.0 58.3 65.9 57.8 79.0 71.0 44.6 72.7 71.6
G-DINO+DH - 64.2 62.4 66.1 65.1 61.3 69.1 57.3 78.2 69.9 45.1 73.2 71.8
ReferDINO (ours) ICCV’25 67.5 65.5 69.6 66.7 62.9 70.7 58.9 80.2 72.3 45.6 74.2 73.1
@@
G-DINO+SH - 66.2 64.2 68.3 65.2 61.6 68.8 58.4 80.9 72.6 45.5 74.2 72.7
G-DINO+DH - 66.7 64.5 69.0 67.6 63.6 71.3 58.7 80.2 71.8 45.3 74.1 73.2
ReferDINO (ours) ICCV’25 69.3 67.0 71.5 68.9 65.1 72.9 61.1 82.1 73.6 46.6 74.2 73.2
Table 1: Comparison on Ref-YouTube-VOS, Ref-DAVIS17, A2D-Sentences and JHMDB-Sentences.
Training. Suppose the ground truth object sequence to be $\boldsymbol{y}=\{\boldsymbol{c}^{t},\boldsymbol{b}^{t},\boldsymbol{m}^{t}\}_{t=1}^{T}$. Following the previous works [4, 3, 23], we apply the Hungarian algorithm for instance matching. Specifically, we select the instance sequence with the lowest matching cost as the positive and assign the remaining sequences as negative. The matching cost is defined as:
This matching cost is computed on individual frames and normalized by the frame number. Here, $\mathcal{L}_{\text{cls}}$ is the focal loss that supervises the binary classification prediction. $\mathcal{L}_{\text{box}}$ sums up the L1 loss and GIoU loss. $\mathcal{L}_{\text{mask}}$ is the combination of DICE loss, binary mask focal loss and projection loss [33]. $\lambda_{\text{cls}}$, $\lambda_{\text{box}}$ and $\lambda_{\text{mask}}$ are scalar weights of individual losses. The model is optimized end-to-end by minimizing the total loss $\mathcal{L}_{\text{total}}$ for positive sequences and only the classification loss $\mathcal{L}_{\text{cls}}$ for negative sequences.
Inference. At inference, we select the best sequence with the highest average classification score as follows:
$\sigma=\underset{i\in[1,N_{s}]}{\arg\max}\frac{1}{T}\sum\nolimits_{t=1}^{T}\boldsymbol{c}^{t}_{i}.$ (6)
5 Experiments
5.1 Experimental Setup
@@
Datasets and Metrics. We evaluate ReferDINO on five public RVOS benchmarks: Ref-YouTube-VOS [31] covers 3,978 videos with 15K descriptions. Ref-DAVIS17 [15] contains 90 videos and 1.5K descriptions. A2D-Sentences [7] contains 3.7K videos and 6.6K descriptions. JHMDB-Sentences [7] contains 928 videos and 928 descriptions. MeViS [6] contains 2K videos and 28K descriptions. On Ref-YouTube-VOS, Ref-DAVIS17 and MeViS, we use region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$ as metrics. For A2D-Sentences and JHMDB-Sentences, we employ mAP, overall IoU (oIoU), and mean IoU (mIoU) metrics.
@@
Protocols. We follow the protocols established in previous works, and all evaluations are conducted using the official code or online platforms. Specifically, the model trained on the training set of Ref-YouTube-VOS is directly evaluated on both Ref-YouTube-VOS and Ref-DAVIS17. Similarly, the model trained on A2D-Sentences is directly evaluated on A2D-Sentences and JHMDB-Sentences. The model is first pretrained on image datasets RefCOCO/+/g [14, 24], and then trained on the RVOS datasets, except for MeViS where the model is trained directly, following [6, 9].
@@
Implementation Details. Our model is built upon the pretrained GroundingDINO, which uses Swin Transformer as the image backbone and BERT as the text backbone. The official source releases two GroundingDINO checkpoints: Swin-T and Swin-B, both of which are covered in our experiments. We freeze the backbones and finetune the cross-modal Transformer with LoRA techniques [11], where the rank is set to $32$. We set $\alpha=0.1$, $L_{t}=3$, and $L_{m}=3$. In the mask decoder, we set the number of sampling points to 16. In the query pruning, we set the drop rate as $50\%$. For MeViS that involves multiple target objects, we follow the practice in [6, 9] to select multiple object trajectories with classification scores higher than a threshold $\sigma=0.3$. For the other datasets, we select the best object trajectory. Extra training details can be found in the supplementary B.
@@
G-DINO-SH - 43.4 39.8 47.1
G-DINO-DH - 45.4 41.9 48.9
ReferDINO (ours) ICCV’25 48.0 43.6 52.3
@@
G-DINO-SH - 45.0 41.3 48.9
G-DINO-DH - 47.9 44.3 51.5
ReferDINO (ours) ICCV’25 49.3 44.7 53.9
@@
5.2 Main Results
Comparisons with SOTA Methods. As shown in Tables 1 and 2, our ReferDINO significantly and consistently outperforms SOTA methods across all five RVOS datasets. Specifically, with the Swin-T backbone, our ReferDINO achieves 48.0% $\mathcal{J}\&\mathcal{F}$ on MeViS, outperforming existing SOTA by 1.6%. On Ref-YouTube-VOS, our ReferDINO achieves 66.4% $\mathcal{J}\&\mathcal{F}$, improving SOTA performance by 2.8%. When a larger backbone Swin-B is applied, ReferDINO further improves $\mathcal{J}\&\mathcal{F}$ to 69.3% on Ref-YouTube-VOS, surpassing SOTA by more than 2.2%. Consistent performance improvements are observed across the other datasets, which demonstrates the superiority of ReferDINO.
Comparisons with Baselines. To further evaluate ReferDINO’s effectiveness, we develop two enhanced variants of GroundingDINO as baselines: G-DINO+SH (static mask head) and G-DINO+DH (dynamic mask head [33]). The former directly dot products the query embedding with the feature map, while the latter uses object-wise feature maps as discussed in Section 4.1. These two mask heads are generic and commonly used in previous RVOS models [3, 40, 23, 25, 45, 9]. Moreover, both baselines are augmented with temporal self-attention modules for cross-frame interaction, as in VideoGroundingDINO [38]. As shown in Tables 1 and  2, although these baselines achieve competitive results on specific datasets (e.g., on Ref-DAVIS17, G-DINO-SH surpasses SOTAs by +0.3% $\mathcal{J}\&\mathcal{F}$ and G-DINO-DH attains +2.7%), they fail to exhibit significant or consistent improvements over existing models across datasets. These results demonstrate that trivial adaptation cannot fully unleash the pretrained capabilities of foundational models. In contrast, our ReferDINO demonstrates robust superiority across different backbones and datasets. The advantages become particularly significant on large benchmarks (Ref-Youtube-VOS and MeViS), where our ReferDINO with Swin-T can even outperform the baselines with Swin-B. More comparison results of GroundingDINO baselines are presented in the supplementary C. These results demonstrate the effectiveness of our approach in adapting foundational models to RVOS task.
@@ 5.3 Ablation Studies
In this section, we mainly analyze the effectiveness of the three proposed components.
Mask Decoder. Our grounding-guided deformable mask decoder utilizes deformable cross-attention (DCA) for grounding-guided adaptive sampling and cross-modal attention (CMA) to further mitigate the impacts of sampling noise. To validate the effectiveness of these two modules, we conduct ablation studies. As shown in Table 3, removing CA and DCA results in performance drops of 0.4% and 2.7% in $\mathcal{J}\&\mathcal{F}$, respectively. These results demonstrate the complementary advantages of the two attention mechanisms, where DCA plays a dominant role in improving mask quality and CA provides further refinement.
        Method         $\mathcal{J}$$&\mathcal{F}$         $\mathcal{J}$         $\mathcal{F}$
        ReferDINO         48.0         43.6         52.3
        w/o CMA         47.6         43.1         52.0
        w/o DCA         45.3         41.8         48.7
Table 3: Ablation study of the mask decoder on MeViS.
Temporal Enhancer. Our object-consistent temporal enhancer consists of a memory-augmented tracker and a cross-modal temporal decoder. We analyze their individual contributions in Table 4. First, while the object queries in GroundingDINO is naturally aligned across frames to some extent, adding a tracker can still improve the temporal consistency, leading to a 0.4% improvement in $\mathcal{J}\&\mathcal{F}$. Second, in the temporal decoder is essential for understanding temporal dynamics and object motion, providing 2.2% improvements in $\mathcal{J}\&\mathcal{F}$. We additionally present visualizations to show the effects of memory-augmented tracker in the supplementary E. These results demonstrates the effectiveness of our object-consistent temporal enhancer.
Figure 5: Qualitative results of our ReferDINO with the SOTA method [23]. ReferDINO performs much better in understanding complex object description involving motion and spatial relationships.
Query Pruning. We explore the impact of purning different numbers of queries. As training with more than 50% of queries on MeViS is highly time-consuming, we conduct this validation on Ref-YouTube-VOS, and present the training costs, inference speed and performance in Table 5. It shows that our query pruning strategy can effectively maintain the performance while significantly reducing the computational costs. For instance, reducing 50% queries with our strategy can dramatically reduce 40.6% FLOPs and 41.2% memory usage, making video-scale training on smaller GPUs possible. Meanwhile, our strategy achieves 51.0 FPS, increasing the inference speed by 10 $\times$ while maintaining consistent performance. In contrast, randomly pruning the same number of queries results in a substantial 29.3% decrease in $\mathcal{J}\&\mathcal{F}$. With our strategy, ReferDINO can be efficiently trained on large RVOS datasets and achieve real-time video processing.
        Method         $\mathcal{J}$$&\mathcal{F}$         $\mathcal{J}$         $\mathcal{F}$
        ReferDINO         48.0         43.6         52.3
@@
Table 4: Ablation study of the temporal enhancer on MeViS.
@@
0 67.6 65.6 69.7 840.3G 24.0G 4.9
15% 67.7 65.7 69.8 672.7G 18.9G 26.5
30% 67.6 65.6 69.6 572.4G 15.9G 47.2
50% 67.5 65.5 69.6 499.3G 14.1G 51.0
Random 50% 38.3 37.1 39.4 499.3G 14.1G 51.0
Table 5: Ablation study of query pruning on Ref-YouTube-VOS. FLOPs and Memory are measured with batch size of $1$ on 6-frame 360P videos, and Pytorch checkpointing disabled. FPS is measured at 360P, following the official evaluation protocol [31].
5.4 Qualitative Analysis
In Figure 5, we provide the qualitative comparisons with the SOTA model [23]. In Figure 6, we present several visualizations of ReferDINO processing various text references within a video. More qualitative results are presented in the supplementary F. These scenarios involve multiple similar objects and complex object descriptions, including rich composite attributions, spatial relationships, and temporal motion. The results clearly demonstrate the effectiveness of ReferDINO in video object segmentation, spatio-temporal reasoning, and vision-language understanding.
6 Conclusion
In this work, we propose ReferDINO for RVOS task to address the challenges in vision-language understanding, dense perception and spatiotemporal reasoning. ReferDINO inherits region-level vision-language alignment from GroundingDINO, and is further endowed with pixel-level object segmentation and cross-modal spatiotemporal reasoning capabilities, through a grounding-guided deformable mask decoder and an object-consistent temporal enhancer. Moreover, we propose a confidence-aware query pruning strategy to accelerate the decoding process. Extensive experiments across five public benchmarks demonstrate that ReferDINO significantly outperforms existing RVOS methods. Moreover, the proposed query pruning strategy significantly improves both training and inference efficiency without compromising performance.
Acknowledgements. This work was supported partially by the NSFC (U22A2095, 6247072922), Guangdong Natural Science Funds Project (2023B1515040025), Guangdong NSF for Distinguished Young Scholar (2022B15-15020009), Guangdong Provincial Key Laboratory of Information Security Technology (2023B1212060026), open research fund of Key Laboratory of Machine Intelligence and System Control, Ministry of Education (No. MISC-202407). We are grateful to the National Supercomputing Center in Guangzhou for providing computing resources.
Supplementary Material
Appendix A Detailed Derivation.
Here, we present the detailed derivation of time complexity for the decoder to demonstrate the effectiveness of our query pruning strategy. GroundingDINO iteratively processes a large number of query embeddings, resulting in significant computational bottlenecks. As GrounddingDINO is based on Deformable DETR [48], its decoder’s computations are mainly occupied by the multi-head self-attention (MHSA) operations and feed-forward layers. MHSA is defined as follows:
$\operatorname{MHSA}(\mathbf{X})=\sum_{i=1}^{H}\mathbf{W}^{o}_{i}\left[\operatorname{softmax}\left(\mathbf{Q}_{i}\mathbf{K}_{i}^{T}\right)\mathbf{V}_{i}\right],$ (7)
$\text{and}\quad\mathbf{Q_{i}}=\mathbf{X}\mathbf{W}_{i}^{q},\quad\mathbf{K}_{i}=\mathbf{X}\mathbf{W}_{i}^{k},\quad\mathbf{V}_{i}=\mathbf{X}\mathbf{W}_{i}^{v},$ (8)
where $\mathbf{X}\in\mathbb{R}^{N\times d}$ indicates the query embeddings, $H$ is the head number, and $\mathbf{W}_{i}^{q}$, $\mathbf{W}_{i}^{k}$, $\mathbf{W}_{i}^{v}$, $\mathbf{W}_{i}^{o}\in\mathbb{R}^{d\times d}$ are the linear weights of $i$-th head. The time complexity of each layer is $O(N^{2}d+Nd^{2})$, hence the total time complexity for a $L$-layer decoder is $O\left(L(Nd^{2}+N^{2}d)\right)$. Then, if we employ the confidence-aware query pruning strategy to identify and retain only $\frac{1}{k}$ queries at each layer, the total time complexity becomes:
$O\left(\left[1+\frac{1}{k^{2}}+\frac{1}{k^{4}}+\cdots+\frac{1}{k^{2(L-1)}}\right]N^{2}d+\left[1+\right.\right.$
$\quad\quad\quad\left.\left.\frac{1}{k}+\frac{1}{k^{2}}+\cdots+\frac{1}{k^{L-1}}\right]Nd^{2}\right)$ (9)
$=$ $O\left(\frac{k^{2}}{k^{2}-1}({\color[rgb]{0.75,0.00,0.00}1-\frac{1}{k^{2L}}})N^{2}d+\frac{k}{k-1}({\color[rgb]{0.75,0.00,0.00}1-\frac{1}{k^{L}}})Nd^{2}\right)$ (10)
$=$ $O\left(\frac{k^{2}}{k^{2}-1}N^{2}d+\frac{k}{k-1}Nd^{2}\right),$ (11)
which is independent of the decoder depth $L$. Note that in equation 10, the items ${\color[rgb]{0.75,0.00,0.00}1-\frac{1}{k^{L}}}$ and ${\color[rgb]{0.75,0.00,0.00}1-\frac{1}{k^{2L}}}$ can be omitted because $\frac{1}{k^{2L}}$ and $\frac{1}{k^{L}}$ are typically much smaller than 1 when $k>1$ and $L$ is a relatively large positive integer.
Appendix B Additional Implementation Details.
FPN. We follow the prior FPN structure [3, 23], which consists of several 2D convolution, GroupNorm and ReLU Layers. Nearest neighbor interpolation is used for upsampling steps.
Implementation Details. The coefficients for losses are set as $\lambda_{cls}=4$, $\lambda_{L1}=5$, $\lambda_{giou}=2$, $\lambda_{dice}=5$, $\lambda_{focal}=5$, $\lambda_{proj}=5$. We apply image augmentation techniques such as flipping, rotation, cropping, and scaling during training, following previous works [40, 25]. The model is pretrained on Ref-COCO/g/+ for 20 epochs with the batch size of 8 for Swin-T and 10 for Swin-B. Then we individually train the models on Ref-YouTube-VOS for 2 epochs and A2D-Sentence for 6 epochs. For MeViS, we train the models for 15 epochs. The number of sampling frames is 6. We use AdamW as the optimizer with the weight decay of 1e-4 and initial learning rate of 5e-5, which is linearly rescaled with the batch size. All experiments are conducted on 8 NVIDIA A800 GPUs.
Appendix C Comparisons with Vanilla Baselines.
As mentioned in Section 2, some previous works [1, 13] also attempted to employ GroundingDINO for RVOS task. Grounded-SAM2 [1] extracts object regions with GroundingDINO and produces masks with SAM2 [28]. Based on this, AL-Ref-SAM2 [13] further incorporates GPT4 [2] to select key frames and boxes. However, such a manner of model ensemble is not end-to-end differentiable, preventing further refinement of RVOS-specific capability. As shown in Table 6, our ReferDINO outperforms these ensemble methods by large margins on MeViS dataset, the largest RVOS benchmark. These results demonstrate the advantages of our end-to-end adaptation approach.
[…]
Video-Swin-T / Swin-T
Grounded-SAM2 [1] 37.4 31.0 43.7 6.1
ReferDINO (ours) 48.0 43.6 52.3 28.0
Video-Swin-B / Swin-B
Grounded-SAM2 [1] 40.5 34.5 46.4 6.1
AL-Ref-SAM 2 [13] 42.8 39.5 46.2 $<$ 6.1
ReferDINO (ours) 49.3 44.7 53.9 26.6
Table 6: Performance comparison on MeViS.
Appendix D Additional Latency Comparisons.
In Table 7, we compare with SOTA dynamic-head methods and our baseline on Ref-Youtube-VOS. The results show that our mask decoder reduces memory usage by 11.2G, enabling superior performance over SOTAs at comparable or faster speed.
[…]
SgMg [25] 62.0 16.5G 50.3
MUTR [42] 64.0 34.0G 41.4
G-DINO+DH 64.2 25.3G 50.2
ReferDINO (Ours) 67.5 14.1G 51.0
Table 7: Comparison of training memory and inference speed.
Appendix E Additional Ablation Studies.
Momentum Coefficient. The momentum coefficient $\alpha$ in our tracker controls the amplitude of memory updating. We visualize a case in Figure 7 to show its impact on temporal consistency. In this case, a smaller $\alpha$ (i.e., long-term memory) yields the best performance. This is because the identification in the initial frames is crucial as it captures the clue “started at the back”. While we set $\alpha=0.1$ by default in main experiments, it is promising to explore an $\alpha$-adaptive strategy in future work. In our main experiments, we set $\alpha=0.1$ by default.
Appendix F More Visualizations.
We provide more visualizations of diverse objects in Figure 8 to demonstrate the robustness of our model.